OpenStudy (anonymous):

HELP FAST PLEASE 10/x + 1/10 = 3/10 HELP FAST PLEASE 10/x + 1/10 = 3/10 @Mathematics

OpenStudy (anonymous):

JUST SOLVE AND SHOW IT BUT JUST GIVE ME THE ASNWER FIRST

OpenStudy (anonymous):

x = 50

OpenStudy (anonymous):

How did u get it explain please!

OpenStudy (anonymous):

i gotchu this is my other sccount

OpenStudy (anonymous):

okay so basically

OpenStudy (anonymous):

...

OpenStudy (anonymous):

just subtract 1/10 from both sides

OpenStudy (anonymous):

just subtract 1/10 frum both sides

OpenStudy (anonymous):

then wt u get 10/x = 2/10 now cross multiply and solve for x

OpenStudy (anonymous):

thanks so much u think u can help me make a simple equation like this for myself?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

so can it be like 3/x+6/3?

OpenStudy (anonymous):

yup but u have to isolate the third term

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

validize to a rational expression

OpenStudy (anonymous):

validize*

OpenStudy (anonymous):

(mulitply both sides by 2)

OpenStudy (anonymous):

doesnt there have to be and equal sign in that equation?

OpenStudy (anonymous):

like there was in my orginal problem?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

can we make one?

OpenStudy (anonymous):

yea sure

OpenStudy (anonymous):

5 bucks

OpenStudy (anonymous):

just the number 5?

OpenStudy (anonymous):

si

OpenStudy (anonymous):

becasue its a a paraphyletic group of organisms that consist of all gill-bearing aquatic vertebrate (or craniate) animals that lacks limbs with digits. Included in this definition are the living hagfish, lampreys, and cartilaginous and bony fish, as well as various extinct related groups. Most fish are "cold-blooded", or ectothermic, allowing their body temperatures to vary as ambient temperatures change, though some of the large active swimmers like white shark and tuna can hold a higher core temperature.[1][2] Fish are abundant in most bodies of water. They can be found in nearly all aquatic environments, from high mountain streams (e.g., char and gudgeon) to the abyssal and even hadal depths of the deepest oceans (e.g., gulpers and anglerfish). At 32,000 species, fish exhibit greater species diversity than any other class of vertebrates.[3] Fish, especially as food, are an important resource worldwide. Commercial and subsistence fishers hunt fish in wild fisheries (see fishing) or farm them in ponds or in cages in the ocean (see aquaculture). They are also caught by recreational fishers, kept as pets, raised by fishkeepers, and exhibited in public aquaria. Fish have had a role in culture through the ages, serving as deities, religious symbols, and as the subjects of art, books and movies. Because the term "fish" is defined negatively, and excludes the tetrapods (i.e., the amphibians, reptiles, birds and mammals) which descend from within the same ancestry, it is paraphyletic, and is not considered a proper grouping in systematic biology. The traditional term pisces (also ichthyes) is considered a typological, but not a phylogenetic classification.

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

are u screwing with me^ lol

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

LML nahh

OpenStudy (anonymous):

its recognized by valid excruations u understanad nah i mean

OpenStudy (anonymous):

okay if u where to solve it? what would it become i needa solve that please

OpenStudy (anonymous):

u tell me i will help u

OpenStudy (anonymous):

how u got a medal the seond u wrote tht 0.o

OpenStudy (anonymous):

sheg gtfo

OpenStudy (anonymous):

thts mah lil bro dnt mind him

OpenStudy (anonymous):

3/x+6/3?=5 ....x=1?

OpenStudy (anonymous):

kk 5 ^

OpenStudy (anonymous):

u too got a medal now gfo

OpenStudy (anonymous):

I GOTTA MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

SHUT UP SHEG BEFORE I TELL MOMMY

OpenStudy (anonymous):

come on is it x=1?

OpenStudy (anonymous):

YUP

OpenStudy (anonymous):

u got it (y)

OpenStudy (anonymous):

nahhhhhh

OpenStudy (anonymous):

wrong as hell flutter feather sleek-feathered one retricehole feather up the fass

OpenStudy (anonymous):

LOLOLOOOOL

OpenStudy (anonymous):

yes x = 1

OpenStudy (anonymous):

byeee guyss!

OpenStudy (anonymous):

bye yasmine muah :*

OpenStudy (anonymous):

yasmeen*

OpenStudy (anonymous):

<3 lol thanks u guys are to funny

OpenStudy (anonymous):

awwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwww

OpenStudy (anonymous):

omg omg omg *faints*